# Positional Encoding in Transformer-Based Time Series Models: A Survey

### v1

@@ Abstract
Recent advancements in transformer-based models have greatly improved time series analysis, providing robust solutions for tasks such as forecasting, anomaly detection, and classification. A crucial element of these models is positional encoding, which allows transformers to capture the intrinsic sequential nature of time series data. This survey systematically examines existing techniques for positional encoding in transformer-based time series models. We investigate a variety of methods, including fixed, learnable, relative, and hybrid approaches, and evaluate their effectiveness in different time series classification tasks. Furthermore, we outline key challenges and suggest potential research directions to enhance positional encoding strategies. By delivering a comprehensive overview and quantitative benchmarking, this survey intends to assist researchers and practitioners in selecting and designing effective positional encoding methods for transformer-based time series models. The source code for the methods and experiments discussed in this survey is available on GitHub [^1] https://github.com/imics-lab/positional-encoding-benchmark.
Introduction
Time series analysis, which involves studying data collected over sequential time intervals, is fundamental across a wide range of domains such as finance, healthcare, and climate science. Traditional approaches, such as Autoregressive Integrated Moving Average (ARIMA) and Seasonal Decomposition of Time Series (STL), have long been the foundation of statistical time series analysis. However, these models are inherently limited in their ability to capture non-linear and long-range dependencies. As a result, machine learning-based approaches, particularly Recurrent Neural Networks (RNNs) and Convolutional Neural Networks (CNNs), have gained popularity due to their ability to model complex temporal dynamics [4, 24].
RNNs, including their more advanced variants like Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU), excel at modeling sequential data by maintaining a hidden state that captures information from previous time steps. These architectures offer several advantages for time series analysis: they naturally handle irregular time intervals and missing data points through their sequential processing, excel at capturing local temporal patterns through their recurrent connections, and exhibit a beneficial ”recency bias” where recent time steps are weighted more heavily than distant ones—a characteristic particularly valuable in applications like financial forecasting and weather prediction. However, RNNs suffer from inherent limitations such as vanishing and exploding gradients, making it difficult to learn dependencies over long time horizons [22].
CNNs, on the other hand, have been adapted for time series by applying convolutional filters over the temporal dimension. While effective at capturing local patterns, their fixed receptive fields limit their ability to model long-term dependencies [20]. Moreover, like RNNs, they may struggle with capturing global patterns and relationships between distant time steps.
In recent years, Transformer-based models, originally developed for natural language processing tasks, have shown remarkable promise in addressing these challenges for time series data [23, 28]. Transformers leverage a self-attention mechanism that enables them to capture both local and global dependencies in data without relying on sequential processing [12]. This architecture allows Transformers to handle long sequences efficiently and mitigates the limitations of recurrence-based models. The self-attention mechanism computes pairwise dependencies between all time steps, providing a global context for each position in the sequence.
However, applying Transformer architectures to time series analysis presents unique challenges. Unlike RNNs, Transformers are inherently permutation-invariant and lack a natural understanding of sequence order [25, 7, 29]. They may also struggle with discontinuous time series data and irregular sampling intervals, necessitating specialized adaptations. Furthermore, although their global attention mechanism excels at capturing long-range dependencies, it can sometimes dilute the model’s ability to focus on significant local patterns. The absence of a natural recency bias, which can be advantageous in many time series applications, may also require explicit architectural modifications.
Positional encoding mechanisms are therefore crucial to enabling Transformer models to distinguish between different time steps in a sequence and to learn temporal relationships effectively [10]. Initial methods, such as the sinusoidal absolute positional encoding introduced by Vaswani et al. [23], used fixed positional information to capture the order of tokens. However, these techniques often struggle with varying sequence lengths and dynamic temporal relationships, which are common in time series data [19, 14].
Recent advancements in positional encoding have introduced learnable and relative [19] positional encodings, which offer more flexibility in capturing temporal dependencies. Hybrid approaches, which combine absolute and relative positional encoding strategies, have also been proposed to leverage the strengths of both methods. These techniques aim to provide a comprehensive representation of positional information, improving the model’s ability to capture complex temporal dependencies while addressing some of the inherent limitations of transformer architectures in time series analysis.
In this survey, we present a systematic analysis of positional encoding methods in transformer-based time series models, categorizing them into three main types: absolute positional encodings, relative positional encodings, and hybrid approaches that combine both techniques. Our methodology encompasses both theoretical analysis and empirical evaluation. We first examine the mathematical foundations and computational characteristics of each encoding method, including their parameter complexity, time complexity, and space requirements. We then evaluate these methods using two distinct transformer architectures: one incorporating batch normalization and another utilizing patch embedding, to assess how architectural choices influence the effectiveness of different positional encoding strategies.
Our experimental evaluation focuses on time series classification tasks and spans eleven diverse datasets, including biomedical signals (EEG, EMG), sensor data, human activity recognition, and financial time series. We prioritize classification over forecasting or anomaly detection because classification tasks provide a more direct measure of a model’s ability to capture temporal dependencies and understand global patterns in time series data. Unlike forecasting, which primarily emphasizes recent temporal context, or anomaly detection, which often relies on local pattern recognition, classification requires a comprehensive understanding of temporal relationships across the entire sequence length. This makes classification particularly well-suited for evaluating positional encoding methods’ effectiveness in capturing both local and global temporal dependencies. This comprehensive evaluation reveals several key findings: advanced methods like Stochastic Positional Encoding (SPE) and Transformer with Untied Positional Encoding (TUPE) [9] consistently outperform traditional approaches, with performance advantages becoming more pronounced for longer sequences. We also find that the effectiveness of positional encoding methods varies significantly with sequence length, data dimensionality, and application domain. These insights provide valuable guidance for practitioners in selecting appropriate positional encoding methods based on their specific application requirements.
Literature Review
Time Series Analysis
Transformer models have emerged as a pivotal advancement in time series analysis, owing to their efficacy in capturing intricate temporal dependencies and long-range interactions. Traditional approaches such as recurrent neural networks (RNNs) [6] and convolutional neural networks (CNNs) [11] have been foundational in time series forecasting, yet they face challenges like vanishing gradients in RNNs and limited ability to model long-term dependencies in CNNs. The Transformer architecture, introduced by Vaswani et al. [23], revolutionized sequence modeling by leveraging self-attention mechanisms. Recent comprehensive surveys by Wen et al. [26] and Li et al. [13] highlight the rapid evolution and current state of transformer applications in time series analysis. This innovation enables Transformers to effectively weigh the relevance of elements within a sequence, irrespective of their temporal distance, thus overcoming the limitations of traditional models. Beyond natural language processing (NLP), Transformers have been adapted for time series applications, demonstrating superior performance in forecasting and representation learning tasks [14, 31]. In time series analysis, handling extensive sequences and complex temporal relationships is pivotal. While traditional models like ARIMA [1] and Exponential Smoothing [8] are constrained by their linear assumptions, transformers have shown remarkable success in capturing nonlinear patterns effectively. Wu et al. [27] introduced Autoformer, incorporating decomposition transformers with auto-correlation for improved long-term series forecasting. Researchers have tailored Transformer architectures to suit the nuances of time series data, leading to substantial enhancements in forecasting accuracy and computational efficiency [29].
Transformer Model Architectures
@@ Introduction
In time series analysis, adaptations of Transformers have led to specialized architectures tailored to exploit their strengths. For instance, the Transformer-based framework for multivariate time series representation learning [29] integrates batch normalization to stabilize training, enhancing model performance on complex multivariate datasets. This framework underscores the significance of normalization techniques in fortifying Transformer models for diverse time series tasks. Moreover, inspired by vision Transformers, patch embeddings have been adapted for time series data. This approach, akin to the Vision Transformer (ViT) paradigm [2], segments time series into patches for localized processing. By encapsulating both local and global patterns, patch embeddings facilitate hierarchical pattern recognition, thereby improving model interpretability and performance in time series analysis.
Positional Encoding Methods
A crucial aspect of employing Transformers in time series tasks is the positional encoding mechanism. Initially proposed by Vaswani et al. [23], fixed positional encodings utilize sinusoidal functions to embed positional information, ensuring the model can differentiate between distinct time steps. Subsequent advancements, such as learnable positional encodings and relative positional encodings introduced by Shaw et al. [19], have further enhanced the model’s ability to capture temporal relationships effectively. Recent innovations in positional encoding methods have further enriched Transformer-based approaches for time series analysis. For instance, Foumani et al. [5] introduced novel techniques like Time Absolute Position Encoding (tAPE) and Efficient Relative Position Encoding (eRPE), aimed at enhancing temporal dependency modeling in Transformer architectures. These methods aim to improve the model’s ability to capture temporal dependencies effectively. Furthermore, the authors developed the ConvTran model, which represents a structural fusion of multiple convolutional layers with a transformer architecture. This hybrid model is specifically designed for the classification of complex multivariate time series data. By combining the strengths of convolutional operations in capturing local patterns with the transformer’s global attention mechanism, ConvTran aims to achieve superior performance in handling diverse and dynamic temporal relationships within time series datasets.
This survey aims to systematically compare various positional encoding methods within these Transformer-based architectures, evaluating their impact on time series forecasting and representation learning. By examining the strengths and weaknesses of each approach, we seek to provide insights into the most effective strategies for incorporating positional information in time series Transformers.
Table 1 provides a comprehensive summary of various positional encoding techniques used in transformer-based time series models. The table categorizes these techniques based on three key criteria: (1) the Technique Type, which can be absolute (Abs) or relative (Rel) positional information; (2) the injection technique, which refers to how the positional information is incorporated into the model, either through additive positional embedding (APE) or manipulating attention matrices (MAM); and (3) whether the positional encoding is learnable during training or fixed. This classification helps in understanding the different approaches and their respective methodologies for embedding temporal information into transformer models.
Table 1: Summary of Positional Encoding Techniques
Method Technique Type Injection Technique Learnable/Fixed Ref.
Sinusoidal Positional Encoding (PE) Absolute Additive Positional Embedding Fixed [23]
Learnable Positional Encoding (LPE) Absolute Additive Positional Embedding Learnable [23]
Relative Positional Encoding (RPE) Relative Manipulating Attention Matrices Fixed [19]
time Absolute Positional Encoding (tAPE) Absolute Additive Positional Embedding Fixed [5]
efficient Relative Positional Encoding (eRPE) Relative Manipulating Attention Matrices Learnable [5]
Transformer with Untied Positional Encoding (TUPE) Rel+Abs Manipulating Attention Matrices Learnable [9]
Convolutional Sinusoidal Positional Encoding (ConvSPE) Relative Manipulating Attention Matrices Learnable [15]
Temporal Positional Encoding (T-PE) Rel+Abs Combined techniques Learnable+Fixed [30]
Table 2: Comparison between the number of parameters, time complexity and space complexity of positional encoding methods.
Method Number of Parameters Time Complexity Sinusoidal PE 0 $\mathcal{O}(Ld)$ Learnable PE $Ld$ $\mathcal{O}(Ld)$ RPE $2(2L-1)dl$ $\mathcal{O}(L^{2}d)$ tAPE $Ld$ $\mathcal{O}(Ld)$ eRPE $(L^{2}+L)l$ $\mathcal{O}(L^{2}d)$ TUPE $2dl$ $\mathcal{O}(Ld+d^{2})$ ConvSPE $3Kdh+dl$ $\mathcal{O}(LKR)$ T-PE $2d^{2}l/h+(2L+2l)d$ $\mathcal{O}(L^{2}d)$
@@
The notations used in Table 2 represent key parameters affecting the computational characteristics of different positional encoding methods. The parameter counts include all learnable parameters across layers, where $l$ represents the number of transformer layers, $h$ is the number of attention heads, and $d$ is the hidden dimension. For methods using relative positioning (RPE, eRPE), we include the full attention matrix parameters. For ConvSPE, K represents the convolutional kernel size that controls the local neighborhood for convolution operations.
Formulation and Analysis of Positional Encoding Methods
@@
This section analyzes positional encoding architectures, examining their mathematical foundations, implementation frameworks, and operational characteristics.
Problem definition and notation

### v2

@@ Abstract
Recent advancements in transformer-based models have greatly improved time series analysis, providing robust solutions for tasks such as forecasting, anomaly detection, and classification. A crucial element of these models is positional encoding, which allows transformers to capture the intrinsic sequential nature of time series data. This survey systematically examines existing techniques for positional encoding in transformer-based time series models. We investigate a variety of methods, including fixed, learnable, relative, and hybrid approaches, and evaluate their effectiveness in different time series classification tasks. Our findings indicate that data characteristics like sequence length, signal complexity, and dimensionality significantly influence method effectiveness. Advanced positional encoding methods exhibit performance gains in terms of prediction accuracy, however, they come at the cost of increased computational complexity. Furthermore, we outline key challenges and suggest potential research directions to enhance positional encoding strategies. By delivering a comprehensive overview and quantitative benchmarking, this survey intends to assist researchers and practitioners in selecting and designing effective positional encoding methods for transformer-based time series models. The source code for the methods and experiments discussed in this survey is available on GitHub.
keywords:
Positional Encoding, Self-Attention, Transformer, Time Series.
1 Introduction
Time series analysis, which involves studying data collected over sequential time intervals, is fundamental across diverse domains including finance, healthcare, environmental monitoring, and industrial systems [1, 2]. The sequential nature of temporal data presents unique modeling challenges that distinguish it from other machine learning tasks, requiring specialized approaches to capture temporal dependencies, seasonal patterns, and long-range relationships effectively.
Traditional statistical methods such as Autoregressive Integrated Moving Average (ARIMA) models have long served as the foundation for time series analysis [1]. While these approaches provide interpretable results and work well for stationary data with clear patterns, they face significant limitations when dealing with non-linear relationships, high-dimensional multivariate data, and complex temporal dynamics that characterize modern time series applications [2].
The emergence of deep learning has revolutionized time series modeling by enabling the capture of complex non-linear patterns and long-range dependencies. Recurrent Neural Networks (RNNs) and their advanced variants, including Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU), became the dominant paradigm for sequential modeling [3, 4]. These architectures excel at processing sequential data through their inherent temporal structure, naturally handling variable-length sequences and maintaining memory of previous time steps. However, RNNs suffer from fundamental limitations including vanishing gradients, sequential processing bottlenecks, and difficulty capturing very long-range dependencies [5].
Convolutional Neural Networks (CNNs) offer an alternative approach to time series modeling by applying convolutional filters across the temporal dimension [6, 7]. While CNNs excel at capturing local temporal patterns and benefit from parallel processing, their fixed receptive fields constrain their ability to model long-term dependencies, and they require deep architectures or dilated convolutions to capture global patterns [8, 9].
The transformer architecture, introduced by Vaswani et al. [10], has fundamentally transformed sequence modeling by replacing recurrent connections with self-attention mechanisms. This paradigm shift enables parallel processing of entire sequences while maintaining the ability to capture long-range dependencies through direct attention computation between all sequence positions [11]. The unprecedented success of transformers in natural language processing, demonstrated by models like BERT [12] and GPT [13], has naturally motivated their application to time series analysis [14, 15, 16] with specialized architectures demonstrating superior performance across forecasting [17, 18, 19], classification [20, 21], and anomaly detection tasks [22, 23].
The adaptation of transformers to time series has revealed both remarkable opportunities and unique challenges. Unlike discrete text tokens with semantic relationships, time series involve continuous values with complex temporal dependencies, irregular sampling patterns, and domain-specific characteristics such as seasonality, trends, and cyclical behavior [24, 25]. The inherent permutation invariance of transformer architectures, while advantageous for certain tasks, becomes problematic for time series where temporal ordering is fundamental to the data’s meaning and predictive power [26].
This challenge has elevated positional encoding from a technical implementation detail to a critical design choice that fundamentally determines model effectiveness. Recent comprehensive analyzes have demonstrated that positional encoding methods significantly impact a model’s ability to handle sequence length extrapolation, temporal pattern recognition, and cross-domain generalization [27, 28]. The unique characteristics of temporal data—including non-stationarity, multi-scale patterns, and irregular sampling—require specialized positional encoding approaches that differ substantially from those developed for natural language processing [29].
Despite extensive research in transformer architectures, the time series community lacks a comprehensive evaluation of positional encoding methods specifically designed for temporal data characteristics. Although there are general transformer surveys [30] and NLP-focused positional encoding reviews are available [27], no systematic study has examined how these methods perform in various time series tasks and architectures. This gap is particularly critical given the unique properties of temporal data, including nonstationarity, seasonality, and varying sampling frequencies [24].
This survey addresses this critical gap by providing the first comprehensive empirical evaluation of positional encoding methods specifically for time series applications. We systematically analyze eight distinct positional encoding approaches across eleven diverse datasets representing different domains, sequence characteristics, and temporal patterns. Our evaluation spans two major transformer architectures to ensure the generalizability of findings and provides practical insights for method selection based on specific application requirements.
Our key contributions include:
  • A systematic comparison of eight positional encoding methods, along with their complexity analysis
  • An evaluation of these methods across diverse time series domains and different transformer architectures
  • The identification of performance patterns related to sequence characteristics such as length and dimensionality
  • practical guidelines for method selection
Results demonstrate that advanced methods like Stochastic Positional Encoding (SPE) and Transformer with Untied Positional Encoding (TUPE) [31] consistently outperform traditional approaches, with effectiveness varying significantly by sequence length, dimensionality, and application domain.
The remainder of this survey is organized as follows: Section 2 provides a comprehensive literature review of transformer-based time series models and positional encoding methods. Section 3 presents background on self attention mechanism positional encoding fundamentals. Section 4 details our systematic analysis of different positional encoding categories, and Sections 5-6 discuss evaluation methodologies and empirical results.
2 Literature Review
2.1 Attention-based models
The evolution of attention mechanisms in time series analysis has progressed through several distinct phases, beginning with early attention-augmented architectures and culminating in sophisticated transformer-based models specifically designed for temporal data. Before full transformer architectures gained prominence, researchers explored attention mechanisms as enhancements to existing sequential models, building upon foundational attention work in neural machine translation [32, 33]. Early work focused on augmenting traditional neural networks with attention capabilities to improve their ability to model temporal dependencies and handle variable-length sequences.
The Dual-Stage Attention-Based RNN (DA-RNN) [34] pioneered the use of attention at both input and temporal dimensions, allowing models to selectively focus on relevant features and time steps. This approach demonstrated significant improvements in financial time series prediction by adaptively extracting relevant dependencies, establishing the foundation for attention-based temporal modeling. Similar dual attention strategies were subsequently developed for multivariate forecasting [35] and clinical time series analysis [36].
Attention-based encoder-decoder architectures adapted sequence-to-sequence learning paradigms for time series forecasting [37]. These models showed that attention mechanisms could effectively handle variable-length sequences and improve long-term prediction accuracy by focusing on relevant historical information. This work demonstrated that attention could address some of the fundamental limitations of purely sequential processing approaches. The success of self-attention motivated the development of specialized variants designed for temporal data characteristics. Cross Attention Stabilized Fully Convolutional Neural Network (CA-SFCN) [38] combined convolutional layers with two types of self-attention: temporal attention (TA) for identifying important time steps and variable attention (VA) for capturing relationships between different variables in multivariate time series. This dual attention approach showed that different attention mechanisms could be specialized for different aspects of temporal modeling. Gated Transformer Networks (GTN) [39] introduced two-tower multi-headed attention architectures that capture discriminative information from input series, using learnable gating mechanisms to balance between different attention patterns. This approach demonstrated that attention mechanisms could be made more adaptive to specific temporal patterns, paving the way for more sophisticated attention-based architectures. The computational challenges of applying attention to long time series motivated the development of efficient attention mechanisms. Sparse attention patterns [40] reduce computational complexity by focusing on subsets of time steps, while maintaining the ability to capture relevant dependencies. These efficiency improvements were crucial for making attention-based approaches practical for real-world time series applications with long sequences. Complementary efficiency strategies include locality-enhanced transformers [11] and reformer architectures [41] that reduce memory bottlenecks in long sequence processing.
Recent developments have focused on adaptive attention mechanisms that adjust based on temporal context [42, 43, 44].
2.2 Transformer-Based Models
Building upon the foundation of attention mechanisms, transformer architectures represent a paradigm shift towards pure attention-based sequence modeling. In time series analysis, adaptations of Transformers have led to specialized architectures tailored to exploit their strengths while addressing temporal-specific challenges. The Transformer-based framework for multivariate time series representation learning [20] integrates batch normalization to stabilize training, enhancing model performance on complex multivariate datasets. This framework underscores the significance of normalization techniques in fortifying Transformer models for diverse time series tasks. The multi-head attention mechanism has been extensively adapted for time series applications, with different heads specializing in various temporal patterns. The success of multi-head attention in transformers inspired specialized variants for temporal data. The Temporal Fusion Transformer (TFT) [45] introduced specialized attention heads for different temporal patterns: some heads focus on short-term dependencies, while others capture long-term trends. The model also incorporates variable selection networks to handle high-dimensional input features common in time series applications. Contemporary approaches have explored hierarchical and multi-scale positional encodings [46, 47, 48].
@@ 2.2 Transformer-Based Models
Moreover, inspired by vision Transformers, patch embeddings have been adapted for time series data. This approach, similar to the Vision Transformer (ViT) paradigm [49], segments time series into patches for localized processing. By encapsulating both local and global patterns, patch embeddings facilitate hierarchical pattern recognition, thereby improving model interpretability and performance in time series analysis.
However, transformers face unique challenges in time series applications. Their inherent permutation invariance necessitates explicit positional information to distinguish temporal ordering [15, 50, 20]. Additionally, the global attention mechanism may dilute focus on local patterns, and the absence of natural recency bias requires architectural adaptations for optimal time series performance.
2.3 Positional Encoding in Time Series
The adaptation of positional encoding methods to time series has revealed fundamental differences between temporal and linguistic data that require specialized approaches.
Positional encoding (PE) mechanisms are therefore crucial for enabling transformers to learn temporal relationships effectively, as self-attention mechanisms are inherently permutation-invariant. [51]. Vaswani et al. [10] initially introduced fixed sinusoidal encodings to embed positional information for temporal differentiation. While these provided a foundational approach, they can struggle with varying sequence lengths and dynamic temporal relationships common in time series data [52, 45]. Subsequently, Shaw et al. [52] advanced the field with learnable and relative positional encodings, enhancing temporal relationship modeling capabilities.
Recent innovations have specifically targeted time series challenges. Foumani et al. [21] developed Time Absolute Position Encoding (tAPE) and Efficient Relative Position Encoding (eRPE) for enhanced temporal dependency modeling, alongside the ConvTran model—a hybrid architecture fusing convolutional layers with transformer mechanisms. Other advanced methods include Transformer with Untied Positional Encoding (TUPE) [31] and Convolutional Stochastic Position Encoding (ConvSPE) [53], which address content-position decoupling and linear complexity maintenance respectively. Hybrid approaches, combining absolute and relative strategies, have also been proposed to leverage complementary strengths for comprehensive temporal information representation, aiming to address some of the inherent limitations of transformer architectures in time series analysis.
Recent work has systematically evaluated PE methods for time series applications [54, 55]. Advanced techniques include frequency-based encodings [56, 57], learnable approaches [58, 59], and hybrid strategies [60, 61]. Specialized methods for irregular time series [62, 63] and long sequences [64, 65] have also been developed.
Table 1: Summary of Positional Encoding Techniques. Abs=Absolute, Rel=Relative, Hyb=Hybrid, Add=Additive, Att=Attention, Comb=Combined, F=Fixed, L=Learnable, M=Mixed.
Method Type Inject. Learn. Params Memory Complex.
Sin. PE [10] Abs Add F 0 $\mathcal{O}(Ld)$ $\mathcal{O}(Ld)$
Learn. PE [10] Abs Add L LdLd $\mathcal{O}(Ld)$ $\mathcal{O}(Ld)$
RPE [52] Rel Att F (2L1)dl(2L{-}1)dl $\mathcal{O}(L^{2}d)$ $\mathcal{O}(L^{2}d)$
tAPE [21] Abs Add F LdLd $\mathcal{O}(Ld)$ $\mathcal{O}(Ld)$
RoPE [66] Hyb Att F 0 $\mathcal{O}(Ld)$ $\mathcal{O}(L^{2}d)$
eRPE [21] Rel Att L 2L12L-1 $\mathcal{O}(L^{2}+L)$ $\mathcal{O}(L^{2})$
TUPE [31] Hyb Att L $2dl$ $\mathcal{O}(Ld{+}d^{2})$ $\mathcal{O}(Ld{+}d^{2})$
ConvSPE [53] Rel Att L 3Kdh+dl3Kdh{+}dl $\mathcal{O}(LKR)$ $\mathcal{O}(LKR)$
T-PE [29] Hyb Comb M $2d^{2}l/h{+}(2L{+}2l)d$ $\mathcal{O}(L^{2}d)$ $\mathcal{O}(L^{2}d)$
ALiBi [67] Rel Att F 0 $\mathcal{O}(L^{2}h)$ $\mathcal{O}(L^{2}h)$
@@ 2.3 Positional Encoding in Time Series
Table 1 provides a comprehensive taxonomy of the positional encoding techniques discussed, based on criteria such as technique type, injection method, learning strategy, parameters, and time complexity. The parameter counts include all learnable parameters across layers, where $l$ represents the number of transformer layers, $h$ is the number of attention heads, and $d$ is the hidden dimension. For methods using relative positioning (RPE, eRPE), we include the full attention matrix parameters. For ConvSPE, K represents the convolutional kernel size that controls the local neighborhood for convolution operations.
3 Positional Encoding Analysis
@@
3.1 Problem definition and notation